A worker's total earnings for one day is $100. He received a $20 fixed payment and consumes 14 hours of leisure. What is the hourly wage rate? Show solutions.


1
Expert's answer
2022-02-13T18:01:54-0500

Let "E" be the worker's earnings, "c" be the fixed pay and "L" be the time consumed for leisure. Also, let "k" be the hourly wage. The relationship between earnings and leisure is given by the equation, "E=c+k(24-L)....(i)".

We are given that, "E=100, c=20,L=14" and we can write the equation "(i)" as, "100=20+k(24-14)". Which we can also write as, "100=20+10k......(ii)". To determine the hourly wage rate, we solve for "k" in equation "(ii)" above.

So,

"100-20=10k\implies 80=10k\implies k=8"

Therefore, the hourly wage rate "k=\$8".